# Spectroscopy of $N=50$ isotones with the valence-space density matrix renormalization group

### v1

@@ 1 Introduction
The first-principles solution of the nuclear many-body problem has witnessed continuous advances to heavier systems and exotic nuclei at the neutron-rich extremes. This success is driven by developments in quantum many-body theory (see, e.g., Refs. Hergert (2020); Hebeler (2021)) and from nuclear forces derived within chiral effective field theory (EFT) Epelbaum et al. (2009); Machleidt and Entem (2011); Hammer et al. (2020). With this it became possible to accurately describe more than one hundred fully interacting nucleons in a controlled way Morris et al. (2018); Arthuis et al. (2020); Hu et al. (2021); Hebeler et al. (2023); Tichai et al. (2024); Arthuis et al. (2024). In medium-mass nuclei, the many-body solution is commonly obtained through expansion methods that admit for mild computational scaling with system size, e.g., using coupled cluster (CC) theory Hagen et al. (2014), the in-medium similarity renormalization group (IMSRG) approach Tsukiyama et al. (2011); Hergert et al. (2016); Heinz et al. (2021); Stroberg et al. (2019), Green’s function theory Dickhoff and Barbieri (2004); Somà et al. (2020), or many-body perturbation theory (MBPT) A. Tichai et al. (2020).
@@
In their basic closed-shell formulation, the many-body expansion is performed around a spherical Hartree-Fock (HF) reference state that is obtained through the solution of the symmetry-restricted HF equations. When targeting nuclei far away from shell closures, including static correlations becomes challenging in this basic expansion. To this end, valence-space methods provide a powerful set of tools to obtain an accurate description of open-shell nuclei. Here the initial Hamiltonian is mapped onto a suitably chosen valence space with only a small number of active nucleons that can be explored through standard diagonalization techniques. In the context of ab initio calculations the construction of effective valence-space operators has been initially performed using MBPT (see, e.g., Ref. Holt et al. (2014) based on modern nuclear forces). More recently, the valence space approach was developed for CC and IMSRG frameworks Tsukiyama et al. (2012); Bogner et al. (2014); Jansen et al. (2014); Stroberg et al. (2017); Sun et al. (2018); Stroberg et al. (2019, 2021). While the final valence-space Hamiltonian diagonalization is technically identical to a shell-model calculation Caurier et al. (2005), this provides a first-principles construction of the valence-space Hamiltonian based on chiral EFT. Still valence-space approaches strongly benefit from the methodological advances for solving large-scale eigenvalue problems Shimizu et al. (2019); Johnson et al. (2018). Alternatively, open-shell nuclei are commonly tackled through multi-reference techniques (see, e.g., Refs. Hergert et al. (2013a); Gebrerufael et al. (2017); Yao et al. (2018); A. Tichai et al. (2018a); Frosini et al. (2022a, b, c)) or symmetry-broken expansions Somà et al. (2013); A. Tichai et al. (2018b); Novario et al. (2020); Demol et al. (2021); Hagen et al. (2022); Tichai et al. (2024).
@@
However, even the most advanced shell-model solvers cannot cope with the dimensions of the Hilbert spaces that emerge in large-scale applications. We recently proposed the merging of the density-matrix renormalization group (DMRG) with the VS-IMSRG as a scalable alternative for the description of low-lying states in strongly correlated nuclei Legeza et al. (2015); Tichai et al. (2023). It was demonstrated that energies can be converged much more rapidly within the VS-DMRG approach, thus, reducing the associated many-body uncertainty. With the resulting VS-DMRG we were further able to successfully reproduce the experimentally observed $N=50$ shell closure in ${}^{78}$i Taniuchi et al. (2019) and link the nuclear shell structure to a drop in total entanglement in the many-body wave-function, revealing an interesting connection between nuclear phenomenology and quantum information (QI) measures.
@@
To put the VS-DMRG approach on par with conventional diagonalization techniques, however, one must be able to describe the same set of nuclear observables. Hence in this work we extend the VS-DMRG approach to electromagnetic transition operators needed for the extraction of $B(E2)$ values. With this at hand, we explore the $N=50$ isotones that are governed by a rapid transition between single-particle and collective excitations Nowacki et al. (2016). Moreover, we show the advantage of the VS-DMRG to capture the underlying physics more efficiently and exemplify this improvement using QI measures.
@@ 2 Methodology
For the VS-DMRG calculations, we employ the VS-IMSRG to calculate from a (free-space) nuclear Hamiltonian $H$ an effective valence-space Hamiltonian. This evolution is mediated by a unitary transformation $U(s)$ controlled by a flow parameter $s$, $H(s)=U^{\dagger}(s)H(0)U(s)$. The VS-IMSRG flow nonperturbatively decouples the valence space to lower and higher lying particle and hole excitations Hergert et al. (2016); Stroberg et al. (2019). Practically, the VS-IMSRG transformation yields a set of second-quantized valence-space operators
@@
where $o_{0}$, $o_{pq}$, and $o_{pqrs}$ denote the normal-ordered zero-, one- and (anti-symmetrized) two-body matrix elements, respectively. The single-particle state $|p\rangle=|nljm_{j}m_{t}\rangle$ gathers all quantum numbers: radial quantum number $n$, orbital angular momentum $l$, total angular momentum $j$ with projection $m_{j}$, and isospin projection $m_{t}$ distinguishing protons and neutrons. The commutator evaluation within the VS-IMSRG transformation naturally induces high-body operators that are truncated at the normal-ordered two-body level, while discarding operators at rank three and beyond Hergert et al. (2016); Heinz et al. (2021). The estimated truncation effect on ground-state observables is of the order of $1-2\%$ Heinz et al. (2021). In practice, we employ the Magnus formulation of the IMSRG which enables solving for the transformation itself instead of the transformed Hamiltonian Morris et al. (2015). We highlight that the IMSRG evolution requires a consistent transformation of other operators as well, e.g., for electromagnetic transition operators. While the leading multipole operators are of one-body character, the commutator evaluation induces non-scalar higher-body components that need to be included explicitly.
Once the VS-IMSRG decoupling is performed, the valence-space Hamiltonian is used as input for a variational density matrix renormalization group (DMRG) calculation White (1992); Schollwöck (2011); Szalay et al. (2015). The fermionic single-particle orbitals in Eq. (1) are mapped onto a one-dimensional tensor topology in occupation-number representation with local dimension $d=2$ (occupied/unoccupied). The length of the chain is given by the number of single-particle states in the valence space $N=\dim\,\mathcal{H}_{1}$. Details of the mapping procedure are discussed in Sec. 4. After the lattice mapping the wave function is represented in a matrix-product-state (MPS) representation. The corresponding wave function of $N$ orbitals is an $N$-dimensional tensor, and the CI coefficient corresponding to a determinant $\boldsymbol{\sigma}=(\sigma_{1},\sigma_{2},\ldots,\sigma_{p},\sigma_{p+1},\ldots,\sigma_{N})$ is expressed as a product of matrices $A_{p}^{\sigma_{p}}$ associated to each orbital $p$ as $|\Psi\rangle=\sum_{\boldsymbol{\sigma}}C_{\boldsymbol{\sigma}}|\boldsymbol{\sigma}\rangle$, where
@@
The size of the MPS tensor would scale exponentially with system size. Therefore, in practice, truncations are employed by setting an upper limit $M$ (the so-called bond dimension) on the number of singular values kept, and hence the maximum size of the MPS factors $A_{p}^{\sigma_{p}}$. The memory cost of storing the largest MPS factors scales as $\sim dM^{2}$. Our DMRG implementation exploits Abelian symmetries $G_{\text{sym}}=U_{N}(1)\,\times\,U_{Z}(1)\,\times\,U_{J_{z}}(1)\,\times\,\mathbb{Z}_{2}$, related to neutron-/proton-number conservation, axial symmetry and parity conservation, thus, yielding block-diagonal MPS tensors with respect to the associated quantum numbers Legeza et al. (2022).
Finally, the MPS components $A_{p}^{\sigma_{p}}$ are variationally optimized in the DMRG algorithm through iterative least-square updates. The tensor space is split according to $\mathcal{H}^{N}=\mathcal{H}^{(\text{left})}\,\otimes\,\mathcal{H}_{p}\,\otimes\,\mathcal{H}_{p+1}\,\otimes\,\mathcal{H}^{(\text{right})}$ where $\mathcal{H}^{(\text{left})}$ ($\mathcal{H}^{(\text{right})}$) denote the left (right) blocks that are formed from precontracted $A$ matrices to the left and right of the sites $p$ and $p+1$, respectively. The DMRG is then executed for a sequence of bond dimensions, and the converged result is obtained by either extrapolating $1/M\rightarrow 0$ or as a function of the truncation error given by the discarded singular values Schollwöck (2005); Legeza and Fáth (1996). In this work, the latter procedure has been used by fixing the truncation error and dynamically adapting the $M$ value above a minimum threshold of $M>1024$ Legeza et al. (2003).
The initial VS-IMSRG decoupling is performed in a single-particle space comprising 15 major harmonic-oscillator shells, i.e., $e_{\text{max}}=\text{max}(2n+l)=14$, and the 3N interaction matrix elements are restricted to $e_{1}+e_{2}+e_{3}\leqslant E_{3\text{max}}=24$ to ensure convergence in heavier systems Miyagi et al. (2022). For all our calculations, we employ the “1.8/2.0 (EM)” NN+3N Hamiltonian from Ref. Hebeler et al. (2011), which is based on chiral EFT nucleon-nucleon (NN) and three-nucleon (3N) interactions. The 3N contributions are approximately taken into account by keeping only up to two-body operators after normal-ordering Hagen et al. (2007); Roth et al. (2012); Hergert et al. (2013b). For our study of $N=50$ isotones we use a $0\hbar\omega$ valence space built on top of a ${}^{60}$a core, i.e., the proton $pf$ shell and neutron $sdg$ shell. The resulting valence space consists of 50 single-particle states (20 protons and 30 neutrons).
@@
Figure 1: Convergence of VS-DMRG and CI energies (top panel) and $B(E2)$ transition strengths (bottom panel) of the lowest $0^{+}$, $2^{+}$, and $4^{+}$ states and their transitions as a function of Hilbert-space dimension for ${}^{80}$n. The largest CI dimension encountered at $T_{\text{max}}=7$ is given by $\dim\mathcal{H}_{A}=10^{9.94}=8.8\times 10^{9}$. The VS-DMRG extrapolated energies and transition strengths (and their uncertainties) are shown by the horizontal bands, see text for details.
@@
Next we study the improved convergence of the DMRG wave function due to its more efficient encoding of quantum correlations in the MPS ansatz compared to the linear CI expansion. For large valence spaces in shell-model calculations, where a full CI solution is not possible, one usually relies on further particle-hole truncations to make the CI diagonalization possible. Therefore, to compare the DMRG truncation in terms of bond dimensions with the particle-hole excitation level $T_{\text{max}}$, we use the dimension of the effective Hilbert space as a common measure to quantify performance. In Fig. 1 we show the ground-state energies of the lowest $0^{+},2^{+}$, and $4^{+}$ states (top panel) and the allowed $B(E2)$ transition strengths (bottom panel) for the case of ${}^{80}$n ($Z=30$). The largest VS-DMRG run was performed at bond dimension $M=10240$, whereas the CI calculations are carried out up to $T_{\text{max}}=7$, i.e., including up to $7$-particle-$7$-hole excitations from the energetically lowest configuration. For all states the VS-DMRG convergence is much faster than for CI and comparable accuracy in energy is obtained at Hilbert space sizes that are reduced by two orders of magnitude. We employ an extrapolation scheme based on taking the largest truncation error $\epsilon$ of the last sweep and using a linear fit to obtained the limit $\epsilon\rightarrow 0$ Legeza and Fáth (1996); Legeza et al. (2003); Wouters and Van Neck (2014); Friesecke et al. (2023). Extrapolation uncertainties are obtained by varying the included data points in the fitting procedures. They are depicted by horizontal bands in Fig. 1.
@@
Similarly, electromagnetic transitions converge more rapidly in the VS-DMRG approach as compared to the CI counterpart. Since the $B(E2)$ values are not variational, convergence is not necessarily monotonic. While for the $2^{+}\rightarrow 0^{+}$ transition the results still reveal a monotonic increase, this is not true for the $4^{+}\rightarrow 2^{+}$ transition. To provide further support of the extrapolation, we have repeated the VS-DMRG calculations for the ${}^{80}$n for a set of fixed truncation errors in the range of $10^{-4}$ and $10^{-5}$. Including the last $3-7$ data points corresponding to the smallest truncation errors (with bond dimensions up to $M\simeq 13000$) we obtain $E_{0^{+}}=-678.06(6)\,\text{MeV}$, $E_{2^{+}}=-676.20(7)\,\text{MeV}$, and $E_{4^{+}}=-675.60(8)\,\text{MeV}$. Similarly, for the electromagnetic transitions we obtains $B(E2;2^{+}\rightarrow 0^{+})=93.5(15)e^{2}\,\text{fm}^{4}$ and $B(E2;4^{+}\rightarrow 2^{+})=49.4(8)e^{2}\,\text{fm}^{4}$. Hence, the extrapolation procedure allows for a robust extraction of $B(E2)$ values in VS-DMRG simulations.
@@ 4 Orbital ordering and DMRG tensor topology
So far the DMRG topology, i.e., the mapping of nuclear orbitals onto a one-dimensional lattice has been left unspecified. In the limit of $M\rightarrow\infty$, the ordering of sites is irrelevant, and all choices will yield the full-space results. However, every simulation at finite bond dimension is sensitive to the employed site ordering. Moreover, for certain topologies the DMRG algorithm may get stuck in local minima. For model Hamiltonians with nearest-neighbour interactions, e.g., a one-dimensional Hubbard Hamiltonian, the ordering problem is less severe since there are no long-range interactions in the system. While the impact of site ordering has been extensively discussed in electronic structure applications (see, e.g., Refs. Chan and Head-Gordon (2002); Legeza and Sólyom (2003); Rissler et al. (2006); Barcza et al. (2011)), there is only a brief documentation for nuclear physics applications Legeza et al. (2015). In general, the problem of finding the “best” ordering is computationally complex, and requires the testing of all $N!$ orbital arrangements in a numerical simulation. For realistic problem sizes with $N\gtrsim 20$, this is clearly impractical, and one relies on heuristic arguments of choosing the DMRG topology. Ideally, such heuristics are guided by our microscopic understanding of the many-body correlations between the interacting particles.
@@
Figure 2: VS-DMRG ground-state energy of ${}^{80}$n as a function of bond dimension obtained for different orbital orderings in the DMRG. The different orderings correspond to: randomly sampled tensor topology with mixed protons and neutrons (purple), energetic shell-model ordering (green), two different BCS orderings (blue/red; see text for details), and quasi-optimal ordering based on single-site entropies (orange).
@@
The efficiency of MPS encoding will be reduced in the presence of long-range correlations along the DMRG chain, i.e., coupling and correlations between distant sites. In this case, the bond dimension has to be significantly increased to capture correlations. As a guiding principle, one must identify the dominant pairwise correlations in the system, and map the associated orbitals next to each other. In nuclear physics such correlations manifest in the superfluid character of open-shell nuclei, i.e., the formation of Cooper pairs between time-reversed states due to short-range attractive interactions. Corresponding entanglement patterns have been observed in pairwise correlation measures in open-shell nuclei Legeza et al. (2015); Robin et al. (2021); Tichai et al. (2023); Pérez-Obiol et al. . In the subsequent analysis, we restrict ourselves to like-particle (pp/nn) pairs, and order states accordingly. We neglect the possibility of the formation of pn pairs that might be important in $N=Z$ nuclei.
@@
Figure 2 shows the VS-DMRG convergence of the ground-state energy of ${}^{80}$n for a selection of DMRG topologies. While all calculations are variationally improved with increasing bond dimension, the individual performance is vastly different. The random ordering has a significant energetic offset compared to the other ones, and converges by far the slowest in this benchmark. In these random arrangements we break the block arrangement of proton/neutron orbits. Following our intuition regarding the relevance of superfluid correlations in nuclei, we next investigate two topologies with time-reversed states mapped next to each other. The first one, labelled “BCS p-/n-” employs a np-block with descreasing $j$ values of the multiplets, i.e.,
@@
whereas the second one, labelled “BCS n+/p-” employs a np-block with increasing neutron orbitals and decreasing proton orbitals, i.e.,
@@
For all investigated bond dimensions, the energetically lowest results were obtained from the “quasi-optimal” ordering (orange) by placing orbitals with largest site entropy around the center of the DMRG chain. The resulting order of multiplets is identical to the one displayed in Eq. (4). However, while the quasi-optimal ordering yields the same order of $j$ multiplets as the BCS n+/p- ordering it differs in the organization of the magnetic substates; the BCS ordering employs an arrangements $m_{j}=\frac{j}{2},-\frac{j}{2},...,\frac{1}{2},-\frac{1}{2}$, whereas in the quasi-optimal one angular-momentum projections are increasing $m_{j}=-\frac{j}{2},...,\frac{j}{2}$. The identical ordering of multiplets explains the closeness of the results already at low bond dimension. In essence, highly entangled orbits are those close to the Fermi surface, and are centered in the chain, while highly excited orbitals are optimally mapped towards the tensor topology’s boundaries. Since it is computationally intractable to exhaust all possible arrangements, we label this ordering as “quasi-optimal”. Clearly, our study shows the high relevance of ordering effects in nuclear applications and demonstrates how the quality of the results is linked to the underlying microscopic structure of a nucleus. Still, with increasing bond dimension, all orderings but the random one yield very similar energies.

### v2

@@ 1 Introduction
The first-principles solution of the nuclear many-body problem has witnessed continuous advances to heavier systems and exotic nuclei at the neutron-rich extremes. This success is driven by developments in quantum many-body theory (see, e.g., Refs. [1, 2]) and from nuclear forces derived within chiral effective field theory (EFT) [3, 4, 5]. With this it became possible to accurately describe more than one hundred fully interacting nucleons in a controlled way [6, 7, 8, 9, 10, 11]. In medium-mass nuclei, the many-body solution is commonly obtained through expansion methods that scale polynomially with system size, e.g., using coupled cluster (CC) theory [12], the in-medium similarity renormalization group (IMSRG) approach [13, 14, 15, 16], Green’s function theory [17, 18], or many-body perturbation theory (MBPT) [19].
@@
In their basic closed-shell formulation, the many-body expansion is performed around a spherical Hartree-Fock (HF) reference state that is obtained through the solution of the symmetry-restricted HF equations. When targeting nuclei far away from shell closures, including static correlations becomes challenging in this basic expansion. To this end, valence-space methods provide a powerful set of tools to obtain an accurate description of open-shell nuclei. Here the initial Hamiltonian is mapped onto a suitably chosen valence space with only a small number of active nucleons that can be explored through standard diagonalization techniques. In the context of ab initio calculations the construction of effective valence-space operators has been initially performed using MBPT (see, e.g., Ref. [20] based on modern nuclear forces). More recently, the valence space approach was developed for CC and IMSRG frameworks [21, 22, 23, 24, 25, 16, 26]. While the final valence-space Hamiltonian diagonalization is technically identical to a shell-model calculation [27], this provides a first-principles construction of the valence-space Hamiltonian based on chiral EFT. Still valence-space approaches strongly benefit from the methodological advances for solving large-scale eigenvalue problems [28, 29]. Alternatively, open-shell nuclei are commonly tackled through multi-reference techniques (see, e.g., Refs. [30, 31, 32, 33, 34, 35, 36]) or symmetry-broken expansions [37, 38, 39, 40, 41, 10].
@@
However, even the most advanced shell-model solvers cannot cope with the dimensions of the Hilbert spaces that emerge in large-scale applications. We recently proposed the merging of the density-matrix renormalization group (DMRG) with the VS-IMSRG as a scalable alternative for the description of low-lying states in strongly correlated nuclei [42, 43]. It was demonstrated that energies can be converged much more rapidly within the VS-DMRG approach, thus, reducing the associated many-body uncertainty. With the resulting VS-DMRG we were further able to successfully reproduce the experimentally observed $N=50$ shell closure in 78Ni [44] and link the nuclear shell structure to a drop in total entanglement in the many-body wave-function, revealing an interesting connection between nuclear phenomenology and quantum information (QI) measures.
@@
To put the VS-DMRG approach on par with conventional diagonalization techniques, however, one must be able to describe the same set of nuclear observables. Hence in this work we extend the VS-DMRG approach to electromagnetic transition operators needed for the extraction of $B(E2)$ values. With this at hand, we explore the $N=50$ isotones that are governed by a rapid transition between single-particle and collective excitations [45, 46]. Moreover, we show the advantage of the VS-DMRG to capture the underlying physics more efficiently and exemplify this improvement using QI measures.
@@ 2 Methodology
For the VS-DMRG calculations, we employ the VS-IMSRG to calculate from a (free-space) nuclear Hamiltonian $H$ an effective valence-space Hamiltonian. This evolution is mediated by a unitary transformation $U(s)$ controlled by a flow parameter $s$, $H(s)=U^{\dagger}(s)H(0)U(s)$. The VS-IMSRG flow nonperturbatively decouples the valence space to lower and higher lying particle and hole excitations [14, 16]. Practically, the VS-IMSRG transformation yields a set of second-quantized valence-space operators
@@
where $o_{0}$, $o_{pq}$, and $o_{pqrs}$ denote the normal-ordered zero-, one- and (anti-symmetrized) two-body matrix elements, respectively. The single-particle state $|p\rangle=|nljm_{j}m_{t}\rangle$ gathers all quantum numbers: radial quantum number $n$, orbital angular momentum $l$, total angular momentum $j$ with projection $m_{j}$, and isospin projection $m_{t}$ distinguishing protons and neutrons. The commutator evaluation within the VS-IMSRG transformation naturally induces high-body operators that are truncated at the normal-ordered two-body level, while discarding operators at rank three and beyond [14, 15]. The estimated truncation effect on ground-state observables is of the order of $1-2\%$ [15]. In practice, we employ the Magnus formulation of the IMSRG which enables solving for the transformation itself instead of the transformed Hamiltonian [47]. We highlight that the IMSRG evolution requires a consistent transformation of other operators as well, e.g., for electromagnetic transition operators. While the leading multipole operators are of one-body character, the commutator evaluation induces non-scalar higher-body components that need to be included explicitly.
Once the VS-IMSRG decoupling is performed, the valence-space Hamiltonian is used as input for a variational density matrix renormalization group (DMRG) calculation [48, 49, 50]. The fermionic single-particle orbitals in Eq. (1) are mapped onto a one-dimensional tensor topology in occupation-number representation with local dimension $d=2$ (labeled by $\sigma_{p}\in\{0,1\}$ referring to unoccupied/occupied orbitals, respectively). The length of the chain is given by the number of single-particle states in the valence space $N=\dim\,\mathcal{H}_{1}$. Details of the mapping procedure are discussed in Sec. 4. After the lattice mapping the wave function is represented in a matrix-product-state (MPS) representation. The corresponding wave function of $N$ orbitals is an $N$-dimensional tensor, and the CI coefficient corresponding to a determinant $\boldsymbol{\sigma}=(\sigma_{1},\sigma_{2},\ldots,\sigma_{p},\sigma_{p+1},\ldots,\sigma_{N})$ is expressed as a product of matrices $A_{p}^{\sigma_{p}}$ associated to each orbital $p$ as $|\Psi\rangle=\sum_{\boldsymbol{\sigma}}C_{\boldsymbol{\sigma}}|\boldsymbol{\sigma}\rangle$, where
@@
The size of the MPS tensor would scale exponentially with system size. Therefore, in practice, truncations are employed by setting an upper limit $M$ (the so-called bond dimension) on the number of singular values kept, and hence the maximum size of the MPS factors $A_{p}^{\sigma_{p}}$. The memory cost of storing the largest MPS factors scales as $\sim dM^{2}$. Our DMRG implementation exploits Abelian symmetries $G_{\text{sym}}=U_{N}(1)\,\times\,U_{Z}(1)\,\times\,U_{J_{z}}(1)\,\times\,\mathbb{Z}_{2}$, related to neutron-/proton-number conservation, axial symmetry and parity conservation, thus, yielding block-diagonal MPS tensors with respect to the associated quantum numbers [51].
Finally, the MPS components $A_{p}^{\sigma_{p}}$ are variationally optimized in the DMRG algorithm through iterative least-square updates. The tensor space is split according to $\mathcal{H}^{N}=\mathcal{H}^{(\text{left})}\,\otimes\,\mathcal{H}_{p}\,\otimes\,\mathcal{H}_{p+1}\,\otimes\,\mathcal{H}^{(\text{right})}$ where $\mathcal{H}^{(\text{left})}$ ($\mathcal{H}^{(\text{right})}$) denote the left (right) blocks that are formed from precontracted $A$ matrices to the left and right of the sites $p$ and $p+1$, respectively. The DMRG is then executed for a sequence of bond dimensions, and the converged result is obtained by either extrapolating $1/M\rightarrow 0$ or as a function of the truncation error given by the discarded singular values [52, 53]. In this work, the latter procedure has been used by fixing the truncation error and dynamically adapting the $M$ value above a minimum threshold of $M>1024$ [54].
The initial VS-IMSRG decoupling is performed in a single-particle space comprising 15 major harmonic-oscillator shells, i.e., $e_{\text{max}}=\text{max}(2n+l)=14$, and the 3N interaction matrix elements are restricted to $e_{1}+e_{2}+e_{3}\leqslant E_{3\text{max}}=24$ to ensure convergence in heavier systems [55]. For all our calculations, we employ the “1.8/2.0 (EM)” NN+3N Hamiltonian from Ref. [56], which is based on chiral EFT nucleon-nucleon (NN) and three-nucleon (3N) interactions. The 3N contributions are approximately taken into account by keeping only up to two-body operators after normal-ordering [57, 58, 59]. For our study of $N=50$ isotones we use a $0\hbar\omega$ valence space built on top of a 60Ca core, i.e., the proton $pf$ shell and neutron $sdg$ shell. The resulting valence space consists of 50 single-particle states (20 protons and 30 neutrons).
@@
Figure 1: Convergence of VS-DMRG and CI energies (top panel) and $B(E2)$ transition strengths (bottom panel) of the lowest $0^{+}$, $2^{+}$, and $4^{+}$ states and their transitions as a function of Hilbert-space dimension for 80Zn. The largest CI dimension encountered at $T_{\text{max}}=7$ is given by $\dim\mathcal{H}_{A}=10^{9.94}=8.8\times 10^{9}$. The VS-DMRG extrapolated energies and transition strengths (and their uncertainties) are shown by the horizontal bands, see text for details.
@@
Next we study the improved convergence of the DMRG wave function due to its more efficient encoding of quantum correlations in the MPS ansatz compared to the linear CI expansion. For large valence spaces in shell-model calculations, where a full CI solution is not possible, one usually relies on further particle-hole truncations to make the CI diagonalization possible. Therefore, to compare the DMRG truncation in terms of bond dimensions with the particle-hole excitation level $T_{\text{max}}$, we use the dimension of the effective Hilbert space as a common measure to quantify performance. In Fig. 1 we show the ground-state energies of the lowest $0^{+},2^{+}$, and $4^{+}$ states (top panel) and the allowed $B(E2)$ transition strengths (bottom panel) for the case of 80Zn ($Z=30$). The largest VS-DMRG run was performed at bond dimension $M=10240$, whereas the CI calculations are carried out up to $T_{\text{max}}=7$, i.e., including up to $7$-particle-$7$-hole excitations from the energetically lowest configuration. For all states the VS-DMRG convergence is much faster than for CI and comparable accuracy in energy is obtained at Hilbert space sizes that are reduced by two orders of magnitude. We employ an extrapolation scheme based on taking the largest truncation error $\epsilon$ of the last sweep and using a linear fit to obtain the limit $\epsilon\rightarrow 0$ [53, 54, 60, 61]. Extrapolation uncertainties are obtained by varying the included data points in the fitting procedures. They are depicted by horizontal bands in Fig. 1.
@@
Similarly, electromagnetic transitions converge more rapidly in the VS-DMRG approach as compared to the CI counterpart. Since the $B(E2)$ values are not variational, convergence is not necessarily monotonic. While for the $2^{+}\rightarrow 0^{+}$ transition the results still reveal a monotonic increase, this is not true for the $4^{+}\rightarrow 2^{+}$ transition. To provide further support of the extrapolation, we have repeated the VS-DMRG calculations for the 80Zn for a set of fixed truncation errors in the range of $10^{-4}$ and $10^{-5}$. Including the last $3-7$ data points corresponding to the smallest truncation errors (with bond dimensions up to $M\simeq 13000$) we obtain $E_{0^{+}}=-678.06(6)\,\text{MeV}$, $E_{2^{+}}=-676.20(7)\,\text{MeV}$, and $E_{4^{+}}=-675.60(8)\,\text{MeV}$. Similarly, for the electromagnetic transitions we obtain $B(E2;2^{+}\rightarrow 0^{+})=93.5(15)e^{2}\,\text{fm}^{4}$ and $B(E2;4^{+}\rightarrow 2^{+})=49.4(8)e^{2}\,\text{fm}^{4}$. Hence, the extrapolation procedure allows for a robust extraction of $B(E2)$ values in VS-DMRG simulations.
@@ 4 Orbital ordering and DMRG tensor topology
So far the DMRG topology, i.e., the mapping of nuclear orbitals onto a one-dimensional lattice has been left unspecified. In the limit of $M\rightarrow\infty$, the ordering of sites is irrelevant, and all choices will yield the full-space results. However, every simulation at finite bond dimension is sensitive to the employed site ordering. Moreover, for certain topologies the DMRG algorithm may get stuck in local minima. For model Hamiltonians with nearest-neighbor interactions, e.g., a one-dimensional Hubbard Hamiltonian, the ordering problem is less severe since there are no long-range interactions in the system. While the impact of site ordering has been extensively discussed in electronic structure applications (see, e.g., Refs. [62, 63, 64, 65]), there is only a brief documentation for nuclear physics applications [42]. In general, the problem of finding the “best” ordering is computationally complex, and requires the testing of all $N!$ orbital arrangements in a numerical simulation. For realistic problem sizes with $N\gtrsim 20$, this is clearly impractical, and one relies on heuristic arguments of choosing the DMRG topology. Ideally, such heuristics are guided by our microscopic understanding of the many-body correlations between the interacting particles.
@@
Figure 2: VS-DMRG ground-state energy of 80Zn as a function of bond dimension obtained for different orbital orderings in the DMRG. The different orderings correspond to: randomly sampled tensor topology with mixed protons and neutrons (purple), energetic shell-model ordering (green), two different BCS orderings (blue/red; see text for details), and quasi-optimal ordering based on single-site entropies (orange).
@@
The efficiency of MPS encoding will be reduced in the presence of long-range correlations along the DMRG chain, i.e., coupling and correlations between distant sites. In this case, the bond dimension has to be significantly increased to capture correlations. As a guiding principle, one must identify the dominant pairwise correlations in the system, and map the associated orbitals next to each other. In nuclear physics such correlations manifest in the superfluid character of open-shell nuclei, i.e., the formation of Cooper pairs between time-reversed states due to short-range attractive interactions. Corresponding entanglement patterns have been observed in pairwise correlation measures in open-shell nuclei [42, 66, 43, 67]. In the subsequent analysis, we restrict ourselves to like-particle (pp/nn) pairs, and order states accordingly. We neglect the possibility of the formation of pn pairs that might be important in $N=Z$ nuclei.
In the DMRG the theoretical uncertainty stems from the following sources: $i)$ the accuracy threshold enforced in the diagonalization of the effective Hamiltonian, $ii)$ the truncation error related to the weight of the discarded renormalized states in the Schmidt spectrum, and $iii)$ the so-called environment error, i.e., the approximate representation of the left or right blocks. The final error is a function of the truncation error [53] once the environmental error is eliminated. Since DMRG is a local optimization procedure, the ordering of orbitals has an important effect on how the environmental error is suppressed and on the required bond dimension to reach a given accuracy. In general, in presence of truncations the DMRG can be trapped in local minima if non-optimal orderings are employed and hence the results are more sensitive to the underlying bond dimension.
@@
To illustrate the importance of orbital ordering, Fig. 2 shows the VS-DMRG convergence of the ground-state energy of 80Zn for a selection of DMRG topologies. While all calculations are variationally improved with increasing bond dimension, the individual performance is vastly different. The random ordering has a significant energetic offset compared to the other ones, converges by far the slowest in this benchmark, and may fluctuate withing the large numerical error induced by the diagonalization and environment error. In these random arrangements we break the block arrangement of proton/neutron orbits. Following our intuition regarding the relevance of superfluid correlations in nuclei, we next investigate two topologies with time-reversed states mapped next to each other. The first one, labeled “BCS p-/n-” employs a np-block with decreasing $j$ values of the multiplets, i.e.,
@@
whereas the second one, labeled “BCS n+/p-” employs a np-block with increasing neutron orbitals and decreasing proton orbitals, i.e.,
@@
For all investigated bond dimensions, the energetically lowest results were obtained from the “quasi-optimal” ordering (orange) by placing orbitals with largest site entropy around the center of the DMRG chain. The resulting order of multiplets is identical to the one displayed in Eq. (4). However, while the quasi-optimal ordering yields the same order of $j$ multiplets as the BCS n+/p- ordering it differs in the organization of the magnetic substates; the BCS ordering employs an arrangement $m_{j}=\frac{j}{2},-\frac{j}{2},...,\frac{1}{2},-\frac{1}{2}$, whereas in the quasi-optimal one angular-momentum projections are increasing $m_{j}=-\frac{j}{2},...,\frac{j}{2}$. The identical ordering of multiplets explains the closeness of the results already at low bond dimension. In essence, highly entangled orbits are those close to the Fermi surface, and are centered in the chain, while highly excited orbitals are optimally mapped towards the tensor topology’s boundaries. Since it is computationally intractable to exhaust all possible arrangements, we label this ordering as “quasi-optimal”. Clearly, our study shows the high relevance of ordering effects in nuclear applications and demonstrates how the quality of the results is linked to the underlying microscopic structure of a nucleus. Still, with increasing bond dimension, all orderings but the random one yield very similar energies.